ENERGY TRANSFER AND LASING OF THE LOWER Ar(3p^{5}4p) STATES IN Ar and He

\underline{JIANDE HAN} and MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

Pulsed lasing from optically pumped noble gas atoms has been demonstrated previously in our laboratory. The lasing relied on a three level scheme, which involved the two lowest energy states of the np^{5}(n+1)p configuration and the lowest energy state of the np^{5}(n+1)s configuration ^{3}P_{2}. Population inversion was achieved using collisional relaxation by helium or argon at pressures near 1 atm. State-to-state energy transfer rate constants are required for modeling these systems. In this study, we have measured the energy transfer and quenching rate constants for the lower Ar(3p^{5}4p) states in collisions with Ar and He. The excited states were populated by pulsed laser excitation of metastable Ar atoms. The metastables were generated by a pulsed electrical discharge. Kinetic processes were interrogated using time-resolved laser induced fluorescence and dispersed fluorescence techniques. For Ar* in collision with Ar, our measured rate constants are in reasonable agreement with previously reported values. For Ar* with He, the rate constants had not been reported previously. We have also investigated the possibility of forming an Ar laser buffered with He or Ar and pumped by a CW diode laser. The lasing experiments will be reported, and the application of the measured rate constants to modeling the lasing system will be described.